# A Deep Dive Into How Open-Source Project Maintainers Review and Resolve Bug Bounty Reports

### v1

@@ 5.2.2 Time commitments and priorities
As expected, bug bounty review is reported as a time-consuming challenge but is not ranked as relatively challenging ($L=10$, $\pi=0.066$). In this context, the benefit of having an improved security posture ($L=33$, $\pi=0.124$) reflects a need to address vulnerabilities regardless of team size and available resources. Further benefits such as detailed reports ($L=8$, $\pi=0.138$) and outsourced work ($L=4$, $\pi=0.064$) help speed up the patching process.
Interviewees expressed struggle in bug bounty review because of time constraints ($I=7$). ”We only have a certain amount of time we can spend on the whole project.” “Outside hobbies” and “family time” were some of the directly stated reasons why it is challenging to resolve reports in a timely manner. In the context of bug bounty review, some interviewees also argued that the actual security impact influences when the report will be prioritized or investigated further based on other tasks already in progress ($I=6$).
All interviewees expressed interest in finding ways to make time for appropriate bug bounty reports with legitimate vulnerabilities ($I=17$). One, in particular, stated how “just having the reports come into us has tightened up some areas of the code where we hadn’t spent much time doing things like audits,” creating time to focus on other issues. Others mention having security in mind while writing code to save time and prevent a particular category of vulnerabilities from being reported ($I=17$), paying close attention to concepts such as “input sanitation” when possible.
@@ 5.2.3 Delegated bounties and support
Although delegated bounties, i.e., bounties for which payments come from and are handled by an entity external to the OSS project under examination, was not the most popular listed benefit and has a lower ranking ($L=9$, $\pi=0.070$) respectively, it is what makes bug bounty support possible for OSS projects without sponsor funding. This observation is also supported by the high number of listings for reducing project costs, which is ranked as the fourth most helpful feature ($L=12$, $\pi=0.121$), and the ability to provide more financial incentives to bug hunters, which is ranked as the third most wanted feature ($L=4$, $\pi=0.145$).
Interviewees emphasized the importance of delegated bounties to bug hunters ($I=4$) because it “makes [them] care more about the issue”. One interviewee stated, “When companies give back to the community, that’s nice, so I think it’s giving a really good incentive.” Others desire additional or more significant financial incentives to get the attention of security researchers and experts to focus on a specific target or write more detailed bug bounty reports in general ($I=2$); thus, encouraging holistic OSS security assessments.
Some also argued the value of having a bug bounty platform administrator to be tagged in discussions and perform additional actions ($I=2$), e.g., opening a pull request for OSS project maintainer contact information via security policy. Other interviewees especially enjoyed the idea of being able to easily get in contact with OSS project maintainers and discuss those potential issues ($I=5$). However, some felt like they were being attacked by hunters ($I=4$), which we present along with the idea of having a “software ego” in Section 5.3.4, and forced to create an account to address unwarranted reports. In particular, one interviewee expressed frustration with the exposure of bug bounty to their project as they argued hunters “snoop around for a bounty in a way that they weren’t before, so I worry that now there’s an expectation [for a payment]”, even if the issue is reported through other means, e.g., email.
@@ 5.3.1 Report quality
Detailed reports is the third least listed benefit and ranked third ($L=8$, $\pi=0.138$) respective to other benefits. Interviewees argued its importance in aiding effective and timely vulnerability resolution ($I=8$), and participants further list how incentivized reporting, e.g., by assigning points, is helpful for hunters to increase their reputation for submitting valid reports and is ranked as the most helpful feature ($L=4$, $\pi=0.167$). However, low-quality or spam reports is the most listed challenge and is ranked as the fifth ($L=22$, $\pi=0.094$) largest issue. Many interviewees felt that this resulted from hunters having a “diverted focus” ($I=10$), focusing on money or CVEs ($L=12$, $\pi=0.141$), which is ranked as the most challenging aspect of bug bounty report review.
@@
As previously mentioned, many interviewees emphasized the importance of receiving detailed reports ($I=10$). In particular, the “most important one is the ease of reproducibility,” which we report in Section 5.4.1. About half of the interviewees ($I=8$) mentioned how detailed reports help with severity assessment, reported in Section 5.3.2, and with bug fixing and testing, reported in Section 5.4.2. “Typically go back and ask, well, how did you do that? What steps did you follow? Can you show me a screenshot or video?” One interviewee argues how especially helpful reports, e.g., having enough detail for easy resolution, should receive additional financial incentives.
@@
Many interviewees describe the challenge of receiving low-quality reports ($I=11$). As one interviewee stated, “understanding what the reporter wants to communicate… I didn’t actually understand what they were saying.” Another interviewee described such reports as “really drunk [and] looks like garbage.” Such reports “pollute” the value of bug bounty reports in general and leave OSS project maintainers in doubt about their legitimacy ($I=2$).
@@
Interviewees argue that the mass amount of low-quality and spam reports is because of a diverted focus on bounties and CVEs ($I=10$). One interviewee told us, “Some of those low-effort reports where you feel like they’re just trying to get the bug bounty money, I think that’s like the biggest annoying part.” One interviewee also argued that some bug hunters only care about receiving a CVE to improve their reputation. On the flip side, another interviewee specifically stated how “nobody is really trying to extort us for money, which is a very common thing that occurs,” reflecting a positive view of bug hunters, who are further described as “very professional” during disclosure.
@@
Some interviewees suspect that a subset of bug hunters they have encountered are at an early stage of their security career ($I=2$), e.g., college students. One interviewee describes that some bug hunters think they are “hot shit researchers, but it was pretty apparent they were quite young… I don’t want to say I was thrilled to be their teacher on this stuff [and] you know, helping them to figure it out.”
@@ 5.3.2 Reported severity
Although not the most listed challenge, overstated severity was ranked fourth ($L=9$, $\pi=0.096$) compared to other challenges during the bug bounty report review process. However, validating impact was the third most listed challenge and ranked fifth ($L=11$, $\pi=0.085$), respectively. Related characteristics include feedback capabilities for report reviewers as a helpful feature ($L=6$, $\pi=0.122$) and a bug hunter reputation system ($L=4$, $\pi=0.145$) as the second most wanted feature.
Several interviewees felt that the severity reported by bug hunters was exact or close enough to the actual severity in a real-world software deployment context ($I=6$), two of which described how hunters only seemed to initially run security tooling or glance at the source code without setting it up before submitting a bug bounty report, and hunters did not “learn the software.” In our interpretation, understanding the source code is a core component of calculating severity because, without it, the context in which the security bug exists may overestimate its actual severity.
A majority of interviewees argued how they received reports where the severity was “overstated” or “inflated” compared to the actual security impact on the software ($I=11$). Further, interviewees noticed that these reports consisted of attack vectors assuming administrative privileges ($I=7$). One interviewee described such attacks as “shooting themselves in the foot” and not a priority since “admin isn’t likely to sabotage their own piece of software”.
In other scenarios, some argued that the bug hunter’s initial severity was almost useless and only accurate in some cases ($I=2$). Stemming from the idea that bug hunters are motivated by bounties, one interviewee explained how reports were “almost guaranteed to be high severity and people would just check off all the maximums to see if they could get that through,” and receive a higher bounty or CVE as a result. Interviewees felt that being able to give feedback on the severity and performing modifications was a beneficial feature ($I=7$). Further, having the ability to review or assign a credibility factor to bug hunters is a wanted feature by OSS project maintainers.
@@ 5.3.3 Duplicate reports
Duplicate reports was the second least listed challenge and ranked sixth ($L=6$, $\pi=0.043$). Though duplicate reports seem not to be a significant issue for OSS project maintainers, receiving them can be “frustrating” and “a waste of time to deal with” ($I=7$).
One interviewee found that the best way to deal with duplicates is to simply ignore them. However, others noted that if they ignored duplicates, it would be more likely that another duplicate would be submitted ($I=2$); therefore, the best action to take is to “acknowledge” the report and mark it as “Informative.” In one instance, an interviewee mentioned how duplicate reports came in waves. They described the experience as “a spree where it’s like certainly four or five people at the same time report the same bug again,” leaving a trail of “useless reports” for OSS project maintainers to review. Interviewees also indicate receiving duplicates of the same issue even after addressing a prior report ($I=4$).
One interviewee also experienced receiving duplicate reports for a legitimate vulnerability, but the actual reports ranged in quality (i.e., some were more detailed than others). They described a scenario that reflects this, “the second user who reported the same problem wrote a very nice issue like with code and screenshots, and whatever tests; but I have to mark valid for the first one because that was the first one… I understood that it was a problem, and marked the second one as a duplicate.” In such cases, the first valid report reviewed would be awarded a bounty, which one interviewee felt was unfair to the bug hunter with the “better” report. Another interviewee describes that “reputation is important because it is the kind of precursor for how trustworthy or capable someone is in the platform, and also it can hurt them.”
@@ 5.3.4 Feelings when receiving reports
Several interviewees brought up the notion of “software ego” when receiving bug bounty reports to review ($I=7$). This term can be described in OSS as an OSS maintainer’s sense of pride, ownership, or attachment to their project. One interviewee argued having a software ego is “definitely a very unhelpful mindset because the fact that you put your code out there for people to use, that commits some responsibility… I’m not gonna say you’re liable if there’s a bug in there, but like at worst, what they report does need to be fixed.” Other interviewees who brought up “software ego” agreed, but one of them still “felt attacked” when a bug bounty report was submitted for their project and had to “not let the ego part of it come into it” ($I=3$).
One interviewee also described how they used to “get excited” or “put everything down” when they first started receiving bug bounty reports. Over time, that mindset shifted to “acknowledge then think.” Five interviewees also mentioned how OSS project maintainers should not “panic” and take their time assessing the bug to realize its impact, then creating a plan of resolution. Another interviewee recommended maintainers be upfront with their intentions and let bug hunters know how long it may take to make judgments about the report to lower expectations of a quick review.
@@ 5.4 Vulnerability reproduction, patching, and management
During bug bounty review, OSS project maintainers argue that being able to reproduce the reported security bug and patching it is critical in ensuring the safety of OSS users. These two stages typically occur before public report disclosure, so any obstacles preventing timely reproducibility or deployment fixes will delay notifying affected users.
@@ 5.4.1 Reproducibility
Interviewees argued that the best-detailed reports included a usable PoC ($I=7$), leading to feasible bug reproduction. One argued how having “the proof of concept makes it a lot easier to actually reproduce the bug… then I don’t have to do the paperwork at the end so it’s all about [patching],” which we present in Section 5.4.2. With a usable PoC, interviewees described how it helps promote a productive workflow and creates additional time to focus on other development aspects ($I=4$), e.g., features. Further, being able to reproduce the vulnerability easily allows OSS project maintainers to understand potential fixes, which we also provide insight for in Section 5.4.2.
On the other hand, interviewees expressed concern with reports that did not have an adequate or existing PoC ($I=5$). Some argue that reports should be required to include a working proof-of-concept ($I=3$), which should also have some way of being verified, e.g., by leveraging large-language models ($I=3$). One interviewee described how bug hunters sometimes “don’t really give clear reproduction steps or they give like a corrupt screenshot so you can’t even see [the bug]… making it harder to actually chase down where the actual issue is.” This is especially important because sometimes it turns out to be a false positive ($I=4$). However, as presented in Section 5.2.2, there is limited time to spend on conducting bug bounty report review, let alone on project management ($I=7$).
@@ 5.4.2 Fixing and testing
Some interviewees described the challenge of determining where in the source code the vulnerability occurs because the bug hunter only provided a PoC and reproduction steps ($I=4$). Another issue interviewees argued is a bug hunter referencing a random line of source code instead of a general area or file where they think the vulnerability may be occurring ($I=2$). One interviewee described this problem with the following, “you know, it’s not hard if you know how this stuff works to search the code base to find a piece of text or a variable name or something and be like, oh, I think this is where the problem is.” This shows that OSS project maintainers benefit from bug hunter guidance since they are more security-oriented and could provide additional insight, even if it is not a bug fix.
Others discussed knowledge gaps with implementing fixes to reported bugs ($I=2$). In one particular instance, an interviewee described a time when they had so much difficulty trying to implement a fix that they “limited the functionality of the product” as a patch to the reported vulnerability, i.e., by removing a feature. Some interviewees mention how communication gaps, e.g., the bug hunter remaining silent or non-responsive, hinder the ability to create a patch ($I=7$) and that it would be “nice to work with [bug hunters] during the [patching] process.”
Finally, some interviewees argued testing is a significant challenge ($I=2$). In one case, the interviewee described how they received fuzzing-oriented bug bounty reports; however, when it came to testing, “for [redacted], this usually happens by feeding junk as input and seeing how [redacted] reacts, so the first problem is getting the gist from the junk and making reproducible test cases out of it.” In other words, even though OSS project maintainers can apply fixes based on bug bounty reports, it can be challenging to create regression test cases for CI/CD purposes ($I=2$).

### v2

@@ 5.2.2 Time commitments and priorities
As expected, bug bounty review is reported as a time-consuming challenge but is not ranked as relatively challenging ($L=10$, $\pi=0.066$). In this context, the benefit of having an improved security posture ($L=33$, $\pi=0.124$) reflects a need to address vulnerabilities regardless of team size and available resources. Further benefits such as detailed reports ($L=8$, $\pi=0.138$) and outsourced work ($L=4$, $\pi=0.064$) help speed up the patching process. Alomar et al. [15] validate this by finding that a broader scope for bounties is likely to incur more time costs and how the lack of detailed information slows down the overall vulnerability resolution process.
Interviewees expressed struggle in bug bounty review because of time constraints ($I=7$). ”We only have a certain amount of time we can spend on the whole project” (P4). “Outside hobbies” and “family time” were some of the directly stated reasons why it is challenging to resolve reports in a timely manner. In the context of bug bounty review, some interviewees also argued that the actual security impact influences when the report will be prioritized or investigated further based on other tasks already in progress ($I=6$).
All interviewees expressed interest in finding ways to make time for bug bounty reports with legitimate vulnerabilities ($I=17$). One, in particular, stated how “just having the reports come into us has tightened up some areas of the code where we hadn’t spent much time doing things like audits” (P8), creating time to focus on other issues. Others mention having security in mind while writing code to save time and prevent a particular category of vulnerabilities from being reported ($I=17$), paying close attention to concepts such as “input sanitation” when possible.
@@ 5.2.3 Delegated bounties and support
Although delegated bounties, i.e., bounties for which payments come from and are handled by an entity external to the OSS project under examination, was not the most popular listed benefit and has a lower ranking ($L=9$, $\pi=0.070$) respectively, it is what makes bug bounty support possible for OSS projects without sponsor funding. This observation is also supported by the high number of listings for reducing project costs, which is ranked as the fourth most helpful feature ($L=12$, $\pi=0.121$), and the ability to provide more financial incentives to bug hunters, which is ranked as the third most wanted feature ($L=4$, $\pi=0.145$). Further, participants from Alomar et al. [15] mentioned that providing extra incentives is not common and that the current bug bounty ecosystem does not incentivize external researchers to go beyond surface-level checks.
Interviewees emphasized the importance of delegated bounties to bug hunters ($I=4$) because it “makes [them] care more about the issue”. One interviewee stated, “When companies give back to the community, that’s nice, so I think it’s giving a really good incentive” (P1). Others desire additional or more significant financial incentives to get the attention of security researchers and experts to focus on a specific target or write more detailed bug bounty reports in general ($I=2$); thus, encouraging holistic OSS security assessments.
Some also argued the value of having a bug bounty platform administrator to be tagged in discussions and perform additional actions ($I=2$), e.g., opening a pull request for OSS project maintainer contact information via security policy. Other interviewees especially enjoyed the idea of being able to easily get in contact with OSS project maintainers and discuss those potential issues ($I=5$). However, some felt like they were being attacked by hunters ($I=4$), which we present along with the idea of having a “software ego” in Section 5.3.4, and forced to create an account to address unwarranted reports. In particular, one interviewee expressed frustration with the exposure of bug bounty to their project as they argued hunters “snoop around for a bounty in a way that they weren’t before, so I worry that now there’s an expectation [for a payment]” (P11), even if the issue is reported through other means, e.g., email.
@@ 5.3.1 Report quality
Detailed reports is the third least listed benefit and ranked third ($L=8$, $\pi=0.138$) respective to other benefits. Interviewees argued its importance in aiding effective and timely vulnerability resolution ($I=8$), and participants further list how incentivized reporting, e.g., by assigning points, is helpful for hunters to increase their reputation for submitting valid reports and is ranked as the most helpful feature ($L=4$, $\pi=0.167$). However, low-quality or spam reports is the most listed challenge and is ranked as the fifth ($L=22$, $\pi=0.094$) largest issue. Alomar et al. [15] also revealed concerns with low-quality or noisy reports from bug bounty hunters. Many interviewees felt that this resulted from hunters having a “diverted focus” ($I=10$), focusing on money or CVEs ($L=12$, $\pi=0.141$), which is ranked as the most challenging aspect of bug bounty report review.
@@
As previously mentioned, many interviewees emphasized the importance of receiving detailed reports ($I=10$). In particular, the “most important one is the ease of reproducibility” (P8), which we report in Section 5.4.1. About half of the interviewees ($I=8$) mentioned how detailed reports help with severity assessment, reported in Section 5.3.2, and with bug fixing and testing, reported in Section 5.4.2. “Typically go back and ask, well, how did you do that? What steps did you follow? Can you show me a screenshot or video?” (P16). One interviewee argues that helpful reports, e.g., having enough detail for easy resolution, should receive additional financial incentives.
@@
Many interviewees describe the challenge of receiving low-quality reports ($I=11$). As one interviewee stated, “understanding what the reporter wants to communicate… I didn’t actually understand what they were saying” (P8). Another interviewee described such reports as “really drunk [and] looks like garbage” (P15). Such reports “pollute” the value of bug bounty reports in general and leave OSS project maintainers in doubt about their legitimacy ($I=2$).
@@
Interviewees argue that the mass amount of low-quality and spam reports is because of a diverted focus on bounties and CVEs ($I=10$). One interviewee told us, “Some of those low-effort reports where you feel like they’re just trying to get the bug bounty money, I think that’s like the biggest annoying part” (P4). One interviewee also argued that some bug hunters only care about receiving a CVE to improve their reputation. On the flip side, another interviewee specifically stated how “nobody is really trying to extort us for money, which is a very common thing that occurs” (P2), reflecting a positive view of bug hunters, who are further described as “very professional” during disclosure.
@@
Some interviewees suspect that a subset of hunters they have encountered are at an early stage of their security career ($I=2$), e.g., college students. One interviewee describes that some bug hunters think they are “hot shit researchers, but it was pretty apparent they were quite young… I don’t want to say I was thrilled to be their teacher on this stuff [and] you know, helping them to figure it out” (P15).
@@ 5.3.2 Reported severity
Although not the most listed challenge, overstated severity was ranked fourth ($L=9$, $\pi=0.096$) compared to other challenges during the bug bounty report review process. However, validating impact was the third most listed challenge and ranked fifth ($L=11$, $\pi=0.085$), respectively. Alomar et al. [15] also provide evidence that CVSS scoring and vulnerability rating mechanisms are insufficient when assessing vulnerability severity, and how misunderstood impact could result in closing reports. Further, Akgul et al.[13] reveal that hunters’ dissatisfaction with responses, e.g., downgraded severities, are particularly challenging to deal with. Related characteristics include feedback capabilities for report reviewers as a helpful feature ($L=6$, $\pi=0.122$) and a bug hunter reputation system ($L=4$, $\pi=0.145$) as the second most wanted feature.
Several interviewees felt that the severity reported by bug hunters was exact or close enough to the actual severity in a real-world software deployment context ($I=6$), two of which described how hunters only seemed to initially run security tooling or glance at the source code without setting it up before submitting a bug bounty report, and hunters did not “learn the software.” On the other hand, Akgul et al. [13] found that hunters feel unclear scopes pose additional challenges, e.g., not knowing what constitutes a valid vulnerability. In our interpretation, understanding the source code is a core component of calculating severity because, without it, the context in which the security bug exists may overestimate its actual severity.
A majority of interviewees argued how they received reports where the severity was “overstated” compared to the actual security impact on the software ($I=11$). Further, interviewees noticed that these reports consisted of attack vectors assuming administrative privileges ($I=7$). One interviewee described such attacks as “shooting themselves in the foot” and not a priority since “admin isn’t likely to sabotage their own piece of software” (P11).
In other scenarios, some argued that the bug hunter’s initial severity was almost useless and only accurate in some cases ($I=2$). Stemming from the idea that bug hunters are motivated by bounties, one interviewee explained how reports were “almost guaranteed to be high severity and people would just check off all the maximums to see if they could get that through” (P2) and receive a higher bounty or CVE as a result. Interviewees felt that being able to give feedback on the severity and performing modifications was a beneficial feature ($I=7$). Further, having the ability to review or assign a credibility factor to bug hunters is a wanted feature by OSS project maintainers.
@@ 5.3.3 Duplicate reports
Duplicate reports was the second least listed challenge and ranked sixth ($L=6$, $\pi=0.043$). Though duplicate reports seem not to be a significant issue for OSS project maintainers, receiving them can be “frustrating” and “a waste of time to deal with” ($I=7$). Akgul et al. [13] found that bug bounty hunters find duplicate reports particularly challenging, showcasing opposing opinions on the value of duplicate reports between these reviewers and hunters.
One interviewee found that the best way to deal with duplicates is to simply ignore them. However, others noted that if they ignored duplicates, it would be more likely that another duplicate would be submitted ($I=2$); therefore, the best action to take is to “acknowledge” the report and mark it as “Informative.” In one instance, an interviewee mentioned how duplicate reports came in waves. They described the experience as “a spree where it’s like certainly four or five people at the same time report the same bug again,” leaving a trail of “useless reports” for OSS maintainers to review (P7). Interviewees also indicate receiving duplicates of the same issue even after addressing a prior report ($I=4$).
One interviewee also experienced receiving duplicate reports for a legitimate vulnerability, but the actual reports ranged in quality (i.e., some were more detailed than others). They described a scenario that reflects this, “the second user who reported the same problem wrote a very nice issue like with code and screenshots, and whatever tests; but I have to mark valid for the first one because that was the first one… I understood that it was a problem, and marked the second one as a duplicate” (P4). In such cases, the first valid report reviewed would be awarded a bounty, which one interviewee felt was unfair to the bug hunter with the “better” report. Another interviewee describes that “reputation is important because it is the kind of precursor for how trustworthy or capable someone is in the platform, and also it can hurt them” (P3).
@@ 5.3.4 Feelings when receiving reports
Several interviewees brought up the notion of “software ego” when receiving bug bounty reports to review ($I=7$). This term can be described in OSS as an OSS maintainer’s sense of pride, ownership, or attachment to their project. One interviewee argued having a software ego is “definitely a very unhelpful mindset because the fact that you put your code out there for people to use, that commits some responsibility… I’m not gonna say you’re liable if there’s a bug in there, but like at worst, what they report does need to be fixed” (P9). Other interviewees who brought up “software ego” agreed, but one of them still “felt attacked” when a bug bounty report was submitted for their project and had to “not let the ego part of it come into it” ($I=3$).
One interviewee also described how they used to “get excited” or “put everything down” when they first started receiving bug bounty reports. Over time, that mindset shifted to “acknowledge then think” (P12). Five interviewees also mentioned how OSS project maintainers should not “panic” and take their time assessing the bug to realize its impact, then creating a plan of resolution. Another interviewee recommended maintainers be upfront with their intentions and let bug hunters know how long it may take to make judgments about the report to lower expectations of a quick review.
@@ 5.4 Vulnerability reproduction, patching, and management
During bug bounty review, OSS project maintainers argue that being able to reproduce the reported security bug and patching it is critical in ensuring the safety of OSS users. These two stages typically occur before public report disclosure, so any obstacles preventing timely reproducibility or deployment fixes will delay notifying affected users. Alomar et al. [15] find that a lack of detailed information in reports for reproducing vulnerabilities hinders informed decision making on how to fix them. Further, Akgul et al. [13] report that bug bounty hunters realize that their PoCs might not be reviewed or fully understood, leading to disagreements in severity, resulting in lower bounty payouts.
@@ 5.4.1 Reproducibility
Interviewees argued that the best-detailed reports included a usable PoC ($I=7$), leading to feasible bug reproduction. One argued how having “the proof of concept makes it a lot easier to actually reproduce the bug… then I don’t have to do the paperwork at the end so it’s all about [patching]” (P5), which we present in Section 5.4.2. With a usable PoC, interviewees described how it helps promote a productive workflow and creates additional time to focus on other development aspects ($I=4$), e.g., features. Further, being able to reproduce the vulnerability easily allows OSS project maintainers to understand potential fixes, which we also provide insight for in Section 5.4.2.
On the other hand, interviewees expressed concern with reports that did not have an adequate or existing PoC ($I=5$). Some argue that reports should be required to include a working proof-of-concept ($I=3$), which should also have some way of being verified, e.g., by leveraging large-language models ($I=3$). One interviewee described how bug hunters sometimes “don’t really give clear reproduction steps or they give like a corrupt screenshot so you can’t even see [the bug]… making it harder to actually chase down where the actual issue is” (P9). This is especially important because sometimes it turns out to be a false positive ($I=4$). However, as presented in Section 5.2.2, there is limited time to spend on conducting bug bounty report review, let alone on project management ($I=7$).
@@ 5.4.2 Fixing and testing
Some interviewees described the challenge of determining where in the source code the vulnerability occurs because the bug hunter only provided a PoC and reproduction steps ($I=4$). Another issue interviewees argued is a bug hunter referencing a random line of source code instead of a general area or file where they think the vulnerability may be occurring ($I=2$). One interviewee described this problem with the following, “you know, it’s not hard if you know how this stuff works to search the code base to find a piece of text or a variable name or something and be like, oh, I think this is where the problem is” (P5). This shows that OSS project maintainers benefit from bug hunter guidance since they are more security-oriented and could provide additional insight, even if it is not a bug fix.
Others discussed knowledge gaps with implementing fixes to reported bugs ($I=2$). In one particular instance, an interviewee described a time when they had so much difficulty trying to implement a fix that they “limited the functionality of the product” (P4) as a patch to the reported vulnerability, i.e., by removing a feature. Some interviewees mention how communication gaps, e.g., the bug hunter remaining silent or non-responsive, hinder the ability to create a patch ($I=7$) and that it would be “nice to work with [bug hunters] during the [patching] process.”
Finally, some interviewees argued testing is a significant challenge ($I=2$). In one case, the interviewee described how they received fuzzing-oriented bug bounty reports; however, when it came to testing, “for [redacted], this usually happens by feeding junk as input and seeing how [redacted] reacts, so the first problem is getting the gist from the junk and making reproducible test cases out of it” (P14). In other words, even though OSS project maintainers can apply fixes based on bug bounty reports, it can be challenging to create regression test cases for CI/CD purposes ($I=2$).